A car that has a mass of 2500kg accelerates to 4 m/s^2. The force in Newton’s that is applied to the car is?
1
Expert's answer
2020-06-22T11:07:39-0400

We can find the force that is applied to the car from the Newton's Second Law of Motion:


"F = ma = 2500 \ kg \cdot 4 \ \dfrac{m}{s^2} = 10000 \ N."

Answer:

"F = 10000 \ N."